Question 19 of 36 Du What co The Kup of Ag SO, can be found here. Number So3IG Next Era gen ®Give Up & View Solution ·Check Answe,。Next 。Previous Y Hint
Answer #1

SO3-2

Ag2SO4 <--> 2Ag+ + SO4-2

Ksp = [Ag]^2[SO4-2]

Ksp = 1.8*10^-4 NOTE: this is very sensitive, please provide Ksp next time so we get the exact value

then

Ksp = [Ag]^2[SO4-2]

1.8*10^-4 = (8.7*10^-3)^2(x)

x = (1.8*10^-4)/((8.7*10^-3)^2) = 2.3781 M